# ZK-APEX: Zero-Knowledge Approximate Personalized Unlearning with EXecutable Proofs

### v1

@@ Abstract
On Vision Transformer (ViT) classification models, our approach recovers approximately 99% Top-1 personalization accuracy while enforcing effective forgetting. We further evaluate the unlearning algorithm on a generative model, OPT125M, trained on the CodeParrot code dataset, achieving $\sim$70% recovery of original accuracy. ZK-SNARK proof generation for the ViT case completes in $\approx$2 hours, which is more than $10^{7}\times$ faster than retraining based verification, with peak memory under 0.7 GB and proof sizes about 400 MB. Together, these results establish the first verifiable personalized unlearning framework practical for deployment on resource constrained edge devices.
Keywords Machine Unlearning $\cdot$ Verifiable Machine Unlearning $\cdot$ ZKML
@@ 1 Introduction
Machine unlearning aims to remove the influence of specific data points from trained models, thereby aligning machine learning systems with privacy regulations such as General Data Protection Regulation (GDPR) mantelero2013eu and California Privacy Rights Act (CPRA) harding2019understanding, which grant individuals the right to be forgotten regulation2016regulation; cao2015towards. Beyond privacy compliance, unlearning is also required for mitigating biases liu2025disentangling, correcting corrupted data, or retracting samples obtained under invalid consent li2025machine. In many practical deployments, such as smartphone keyboards hard2018federated; singhal2021federated, photo categorization models (for example Google Photos and Apple Photos) gunter2024apple, and voice assistants du2024communication, the model provider, typically a company, distributes a pretrained model to edge client devices. Each device personalizes the model on locally collected private data, as recently demonstrated in Apple Intelligence gunter2024apple, which enables large-scale on-device adaptation through lightweight foundation models. When the provider later requests the removal of certain samples, it must verify that each client has executed the unlearning correctly, while the clients must not expose their personalized models, which contain sensitive private information balle2022reconstructing; leino2020stolen. This creates a verification challenge: the client, who holds a locally personalized model, must convince the model provider that the specified data has been erased, without exposing the underlying model parameters or private samples. The problem is further complicated in edge environments, where storage, compute, and communication constraints make full retraining or model sharing infeasible. Consequently, there is a growing need for privacy-preserving verification mechanisms that can attest to correct unlearning on-device. A natural direction for enabling such verifiable unlearning is the use of zero-knowledge succinct non-interactive arguments of knowledge (ZK-SNARKs), a cryptographic primitive that enables verifiable computation without revealing private information. ZK-SNARKs allow a prover (e.g., an edge device) to generate a compact proof that a computation, such as an unlearning update, has been performed correctly, while the verifier (e.g., the model provider) can efficiently check its validity without accessing the personalized model parameters or intermediate values kang2022scaling; maheri2025telesparse; rabanser2025confidential.
@@
However, performing unlearning in personalized settings on edge devices introduces distinctive challenges. Unlike centralized unlearning frameworks kurmanji2023towards; thudi2022unrolling; chundawat2023can; chundawat2023zero; jia2023model, where the model provider can compute and distribute global unlearning updates, directly applying such updates to locally personalized models often degrades their task-specific adaptation. Moreover, naive strategies that require the provider to send the forget-set to each client and demand proof of its removal are infeasible for several reasons. First, generating zero-knowledge proofs (ZKP) for complete unlearning procedures such as multi-epoch gradient ascent graves2021amnesiac; thudi2022unrolling, retraining bourtoule2021machine; eisenhofer2025verifiable; yu2023split; xia2025edge, or fine-tuning over the retained dataset hong2024dissecting imposes significant computational and memory overhead qu2025zkgpt; maheri2025telesparse; sun2024zkdl. Compared to inference, producing ZKPs for training is substantially more complex qu2025zkgpt; sun2024zkdl. Prior studies abbaszadeh2024zero; garg2023experimenting; waiwitlikhit2024trustless have explored proofs of training for different machine learning models. However, these methods remain prohibitively expensive and impractical for real-world deployment, especially on edge devices. Second, sharing the raw forget-set with clients is undesirable, as these samples are typically sensitive in unlearning applications xia2025edge; nguyen2025survey. These limitations motivate a new paradigm for verifiable personalized unlearning that is both computationally tractable and privacy-preserving. The key objective is to develop an unlearning mechanism that (i) effectively removes the influence of designated samples from personalized models without sacrificing their local performance, and (ii) remains compatible with ZKP, ensuring that the verification of unlearning is computationally tractable for edge devices. In this work, we address this challenge by introducing a ZK-friendly approximate unlearning framework tailored for personalized models in resource-constrained edge environments.
@@
Recent studies have shown that information in deep networks tends to localize within specific neurons or filters ghorbani2020neuron; lin2020hrank. This observation suggests that removing or masking a small subset of highly influential weights can effectively erase the knowledge associated with targeted samples. Building on this insight, we design our unlearning request as a saliency-based pruning operation, where each weight is scored according to its contribution to the forget-set, using first- and second-order statistics such as gradients and curvature. Masking the weights with the highest scores removes the influence of the forget samples from the model jia2023model; hong2024dissecting. However, directly applying such a mask to a personalized model substantially degrades its accuracy, since personalization relies on locally adapted features that often overlap with those associated with the forget-set. To mitigate this, we formulate a compensation step grounded in the Optimal Brain Surgeon (OBS) framework lecun1989optimal; kurtic2022optimal, which computes a second-order weight adjustment that restores performance on the personalized data while maintaining high loss on the forget domain. This two-part procedure achieves the first goal of effective unlearning on personalized models while preserving their utility. The second goal is verifiability under zero knowledge (ZK). Our algorithm is inherently ZK-friendly because it is zero-shot, performing no stochastic training or iterative optimization after applying the unlearning transformation, and thus avoids the randomness of SGD, which has been shown to enable forging attacks in verifiable unlearning DBLP:conf/icml/ZhangCSL24. Instead, the client only needs to prove that (i) the specified mask correctly zeroed the targeted parameters and (ii) the compensation weights were computed according to the prescribed rule from committed inputs, including the Fisher information matrix of the personalized model (evaluated on the personalization set) and the public mask. These operations reduce to sparse matrix–vector computations that can be efficiently verified within a ZK-SNARK circuit. As a result, we introduce ZK-APEX, Zero-Knowledge Approximate Personalized Unlearning with EXecutable Proofs, which enables practical and privacy-preserving verification of personalized unlearning across diverse deployment scenarios. ZK-APEX achieves proof generation that is even more computationally efficient than typical ZK-SNARK-based inference verification on a single sample.
@@ 2 Related Works and Background
Machine unlearning was first introduced to ensure that trained models could forget specific data as if it had never been used cao2015towards; bourtoule2021machine; graves2021amnesiac. Approaches such as SISA bourtoule2021machine and CAUSE xia2025edge achieve this through retraining on data partitions or maintaining multiple checkpoints of the training process. While these methods guarantee deletion in theory, they depend on access to the full dataset and impose substantial storage and computation costs, making them impractical for large-scale or distributed deployments.
To address these limitations, approximate unlearning methods modify the trained model directly to emulate the effect of retraining kurmanji2023towards; golatkar2020forgetting; chundawat2023can; thudi2022unrolling. A common formulation is class unlearning, where the goal is to erase representations associated with particular semantic categories, while preserving the model’s ability to generalize to the remaining classes chundawat2023zero; seo2025revisiting; fan2024challenging. The main challenge lies in suppressing class-specific information without damaging the broader structure of shared features, as excessive removal can harm retained-class accuracy.
More recently, studies have shown that neural representations are spatially and structurally localized frantar2023sparsegpt; meng2022locating, motivating sparsity-based unlearning approaches that identify and mask a small set of influential weights to remove targeted information fan2023salun; jia2023model; pochinkov2024dissecting. Such methods have been successfully applied to both discriminative architectures jia2023model and generative models fan2023salun; pochinkov2024dissecting such as large language models, demonstrating that selective parameter masking can effectively achieve forgetting, while preserving global performance. Despite their efficiency, existing approximate unlearning approaches generally assume a centralized setup, where the model provider performs unlearning on a shared global model using the forget-set data. In contrast, our work considers unlearning in personalized models that have been locally adapted using private client data, aiming to remove provider-specified class information while retaining personalization-specific knowledge.
@@
ZKPs goldreich1994definitions enable verifiable ML by allowing a prover to demonstrate correct computation without revealing private inputs or intermediate states. ZK-SNARKs ITCS:BCCT12; kilian1992note; micali2000computationally are widely used for ML verification owing to their short proofs, fast verification, and modular circuits that can be partitioned across layers maheri2025telesparse; liu2021zkcnn; lee2024vcnn; sun2024zkllm; zcash-halo2. While suitable for large-scale auditing, proof generation remains costly in time and memory kang2022scaling.
Verification of machine unlearning ensures that models have genuinely forgotten target data. Existing methods include backdoor and sensitivity-based verification, which use poisoned or sensitivity of samples to test deletion sommer2022athena; gao2024verifi; guo2023verifying; zhou2025truvrf, but can be bypassed by a dishonest prover, and reproducing verification, which replays unlearning traces via proofs of learning thudi2022necessity; weng2024proof; eisenhofer2025verifiable but remains computationally heavy. Recent cryptographic frameworks instantiate proofs of unlearning with SNARKs and hash chains eisenhofer2025verifiable, yet these verify exact unlearning (full retraining), making proof generation impractical. Furthermore, they are vulnerable to the stochasticity of SGD DBLP:conf/icml/ZhangCSL24: an adversary can exploit minibatch randomness to retrain on retained samples whose gradients mimic those of the removed data, producing an apparently valid proof even though the model still encodes the forget-set information. This motivates ZK-friendly approximate unlearning methods that achieve verifiability without retraining.
@@ 2.1 Optimal Brain Surgeon
Second-order pruning methods such as Optimal Brain Damage lecun1989optimal and Optimal Brain Surgeon hassibi1993optimal introduced a principled approach to compensate for parameter removal by modeling loss curvature. By expanding the loss near an optimum, these methods derive closed-form updates that minimally affect performance after zeroing selected weights. Later work proposed scalable variants using empirical Fisher information amari1998natural, block-diagonal curvature singh2020woodfisher, and structured sparsity kurtic2022optimal, enabling efficient use in pruning, quantization, and compression of modern networks dong2017learning; wang2019eigendamage; kuznedelev2024cap.
@@ 3 System and Problem Formulation
A client holds a private personalization set $D_{p}$ and obtains a personalized model by applying a fixed operator $P$ (e.g., short-horizon SGD, low-rank adapters such as LoRA hu2022lora, or linear probes): $\theta_{p}\;=\;P(\theta_{0};\,D_{p}).$
@@
Parties and data. The model provider holds $\theta_{0}$, the pretraining corpus $D$, the forget/retain split $(D_{f},D_{r})$, and publishes traceability artifacts $\Psi$ (e.g., a sparse mask $m^{\star}$). The client holds private data $D_{p}$ and its personalized model $\theta_{p}=P(\theta_{0};D_{p})$, and executes $U(\cdot;\Psi)$ locally to obtain $\theta_{u}$. As in prior work kang2022scaling; weng2023pvcnn; maheri2025telesparse, the model architecture is public; model weights ($\theta_{p},\theta_{u}$) remain private.
@@
Proof system primitives. In a ZK setting, verification is conducted through an arithmetic circuit that encodes the agreed computation $U$. A concise overview of circuit representations, constraint systems, and polynomial commitment schemes used in ZK-SNARKs is provided in Appendix B. Here, the client acts as the prover and generates a proof $\pi$ attesting that the committed personalized model $\theta_{p}$ was correctly transformed into $\theta_{u}$ under the public artifact $\Psi$. The provider acts as the verifier and checks the validity of $\pi$ without accessing any private inputs. The public inputs to the circuit are the traceability artifact $\Psi$ and the cryptographic commitments $\mathsf{Com}(\theta_{p})$ and $\mathsf{Com}(\theta_{u})$, which bind model parameters to their corresponding proofs without revealing them. Formal definitions of the proof system properties—soundness, correctness, and zero-knowledge—as well as the binding and hiding guarantees of the commitment scheme are provided in Appendix C. We assume that a valid commitment $\mathsf{Com}(\theta_{p})$ has been established beforehand, for example through a proof-of-training or proof-of-personalization protocol garg2023experimenting; abbaszadeh2024zero; sun2024zkdl.
@@
Threat model. Adversaries are computationally bounded. (i) Privacy: The proof is zero knowledge (see Appendix B); it reveals no information about $\theta_{p}$, $\theta_{u}$, or $D_{p}$ beyond the public inputs $(\Psi,\mathsf{Com}(\theta_{p}),\mathsf{Com}(\theta_{u}))$. The model architecture is public, whereas all model weights and client-specific statistics remain part of the private witness. (ii) Integrity: By knowledge soundness, the verifier accepts only if the publicly committed $\theta_{u}$ equals the (deterministic) output of the agreed operator $U$ applied to the committed $\theta_{p}$ under $\Psi$. After acceptance, all subsequent predictions and updates must reference $\mathsf{Com}(\theta_{u})$, ensuring the deployed model is exactly the certified output of $U$. Operationally, this means the next predictions are produced by the committed $\theta_{u}$, not by any model that still encodes forget-set patterns. All security properties are inherited from Halo2 zcash-halo2; formal definitions and assumptions appear in Appendix C.
@@
Scope. We do not address secure erasure of the client’s pre-unlearning model, as an archival copy of $\theta_{p}$ may persist on the device. Secure deletion is a challenging problem in its own right, often requiring trusted hardware-based method hunt2018ryoan; wu2024secgpt, and is considered orthogonal to our goal of verifiable unlearning. Our guarantee is verifiable use: after acceptance, any future inference or update must reference $\mathsf{Com}(\theta_{u})$, so deployed predictions are produced by the certified unlearned model, and do not rely on parameters that encode forget-set patterns. Operational controls (e.g., requiring a valid proof per update/inference) can enforce this usage constraint, but the physical deletion of prior weights is out of scope.
Empirical evaluation of unlearning quality is conducted independently of these roles; zk certification concerns only procedural correctness, whereas alignment in (3) pertains to the machine-learning evaluation of unlearning quality.
@@
Design goals for proof efficiency. We aim for ZK circuits that (i) minimize the number of constraints by verifying operator-specific identities instead of re-running heavy optimization (e.g., SGD) or training inside the circuit; (ii) utilise only linear operations (matrix–vector products, inner products), to avoid non-linear activations and large lookup tables that dominate SNARK cost maheri2025telesparse; and (iii) avoid stochastic optimization inside the circuit — we verify deterministic operator identities and eliminate randomness  — thereby preventing prover adaptivity and the forging attacks observed in verification–unlearning settings that leverage SGD randomness DBLP:conf/icml/ZhangCSL24.

### v2

@@ Abstract
On Vision Transformer (ViT) classification models, our approach recovers approximately 99% Top-1 personalization accuracy while enforcing effective forgetting. We further evaluate the unlearning algorithm on a generative model, the autoregressive language model OPT125M, trained on the CodeParrot code dataset, achieving $\sim$70% recovery of original accuracy. ZK-SNARK proof generation for the ViT case completes in $\approx$2 hours, which is more than $10^{7}\times$ faster than retraining based verification, with peak memory under 0.7 GB and proof sizes about 400 MB. Together, these results establish the first verifiable personalized unlearning framework practical for deployment on resource constrained edge devices.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
@@ 1 Introduction
Machine unlearning aims to remove the influence of specific data points from trained models, thereby aligning machine learning systems with privacy regulations such as General Data Protection Regulation (GDPR) Mantelero (2013) and California Privacy Rights Act (CPRA) Harding et al. (2019), which grant individuals the right to be forgotten Regulation (2016); Cao and Yang (2015). Beyond privacy compliance, unlearning is also required for mitigating biases Liu et al. (2025), correcting corrupted data, or retracting samples obtained under invalid consent Li et al. (2025). In many practical deployments, such as smartphone keyboards Hard et al. (2018); Singhal et al. (2021), photo categorization models (for example Google Photos and Apple Photos) Gunter et al. (2024), and voice assistants Du et al. (2024), the model provider, typically a company, distributes a pretrained model to edge client devices. Each device personalizes the model on locally collected private data, as recently demonstrated in Apple Intelligence Gunter et al. (2024), which enables large-scale on-device adaptation through lightweight foundation models. When the provider later requests the removal of certain samples, it must verify that each client has executed the unlearning correctly, while the clients must not expose their personalized models, which contain sensitive private information Balle et al. (2022); Leino and Fredrikson (2020). This creates a verification challenge: the client, who holds a locally personalized model, must convince the model provider that the specified data has been erased, without exposing the underlying model parameters or private samples. The problem is further complicated in edge environments, where storage, compute, and communication constraints make full retraining or model sharing infeasible. Consequently, there is a growing need for privacy-preserving verification mechanisms that can attest to correct unlearning on-device. A natural direction for enabling such verifiable unlearning is the use of zero-knowledge succinct non-interactive arguments of knowledge (ZK-SNARKs), a cryptographic primitive that enables verifiable computation without revealing private information. ZK-SNARKs allow a prover (e.g., an edge device) to generate a compact proof that a computation, such as an unlearning update, has been performed correctly, while the verifier (e.g., the model provider) can efficiently check its validity without accessing the personalized model parameters or intermediate values Kang et al. (2022); Maheri et al. (2025); Rabanser et al. (2025).
@@
However, performing unlearning in personalized settings on edge devices introduces distinctive challenges. Unlike centralized unlearning frameworks Kurmanji et al. (2023); Thudi et al. (2022a); Chundawat et al. (2023a; b); Jia et al. (2023), where the model provider can compute and distribute global unlearning updates, directly applying such updates to locally personalized models often degrades their task-specific adaptation. Moreover, naive strategies that require the provider to send the forget-set to each client and demand proof of its removal are infeasible for several reasons. First, generating zero-knowledge proofs (ZKP) for complete unlearning procedures such as multi-epoch gradient ascent Graves et al. (2021); Thudi et al. (2022a), retraining Bourtoule et al. (2021); Eisenhofer et al. (2025); Yu et al. (2023); Xia et al. (2025), or fine-tuning over the retained dataset Hong et al. (2024) imposes significant computational and memory overhead Qu et al. (2025); Maheri et al. (2025); Sun et al. (2024a). Compared to inference, producing ZKPs for training is substantially more complex Qu et al. (2025); Sun et al. (2024a). Prior studies Abbaszadeh et al. (2024); Garg et al. (2023); Waiwitlikhit et al. (2024) have explored proofs of training for different machine learning models. However, these methods remain prohibitively expensive and impractical for real-world deployment, especially on edge devices. Second, sharing the raw forget-set with clients is undesirable, as these samples are typically sensitive in unlearning applications Xia et al. (2025); Nguyen et al. (2025). These limitations motivate a new paradigm for verifiable personalized unlearning that is both computationally tractable and privacy-preserving. The key objective is to develop an unlearning mechanism that (i) effectively removes the influence of designated samples from personalized models without sacrificing their local performance, and (ii) remains compatible with ZKP, ensuring that the verification of unlearning is computationally tractable for edge devices. In this work, we address this challenge by introducing a ZK-friendly approximate unlearning framework tailored for personalized models in resource-constrained edge environments.
@@
Recent studies have shown that information in deep networks tends to localize within specific neurons or filters Ghorbani and Zou (2020); Lin et al. (2020). This observation suggests that removing or masking a small subset of highly influential weights can effectively erase the knowledge associated with targeted samples. Building on this insight, we design our unlearning request as a saliency-based pruning operation, where each weight is scored according to its contribution to the forget-set, using first- and second-order statistics such as gradients and curvature. Masking the weights with the highest scores removes the influence of the forget samples from the model Jia et al. (2023); Hong et al. (2024). However, directly applying such a mask to a personalized model substantially degrades its accuracy, since personalization relies on locally adapted features that often overlap with those associated with the forget-set. To mitigate this, we formulate a compensation step grounded in the Optimal Brain Surgeon (OBS) framework LeCun et al. (1989); Kurtic et al. (2022), which computes a second-order weight adjustment that restores performance on the personalized data while maintaining high loss on the forget domain. This two-part procedure achieves the first goal of effective unlearning on personalized models while preserving their utility. The second goal is verifiability under zero knowledge (ZK). Our algorithm is inherently ZK-friendly because it is zero-shot, performing no stochastic training or iterative optimization after applying the unlearning transformation, and thus avoids the randomness of SGD, which has been shown to enable forging attacks in verifiable unlearning Zhang et al. (2024). Instead, the client only needs to prove that (i) the specified mask correctly zeroed the targeted parameters and (ii) the compensation weights were computed according to the prescribed rule from committed inputs, including the Fisher information matrix of the personalized model (evaluated on the personalization set) and the public mask. These operations reduce to sparse matrix–vector computations that can be efficiently verified within a ZK-SNARK circuit. As a result, we introduce ZK-APEX, Zero-Knowledge Approximate Personalized Unlearning with EXecutable Proofs, which enables practical and privacy-preserving verification of personalized unlearning across diverse deployment scenarios. ZK-APEX achieves proof generation that is even more computationally efficient than typical ZK-SNARK-based inference verification on a single sample.
@@
Our open-source implementation is available at https://github.com/mammadmaheri7/ZK-APEX.
@@ 2 Related Works and Background
Machine unlearning was first introduced to ensure that trained models could forget specific data as if it had never been used Cao and Yang (2015); Bourtoule et al. (2021); Graves et al. (2021). Approaches such as SISA Bourtoule et al. (2021) and CAUSE Xia et al. (2025) achieve this through retraining on data partitions or maintaining multiple checkpoints of the training process. While these methods guarantee deletion in theory, they depend on access to the full dataset and impose substantial storage and computation costs, making them impractical for large-scale or distributed deployments.
To address these limitations, approximate unlearning methods modify the trained model directly to emulate the effect of retraining Kurmanji et al. (2023); Golatkar et al. (2020); Chundawat et al. (2023a); Thudi et al. (2022a). A common formulation is class unlearning, where the goal is to erase representations associated with particular semantic categories, while preserving the model’s ability to generalize to the remaining classes Chundawat et al. (2023b); Seo et al. (2025); Fan et al. (2024). The main challenge lies in suppressing class-specific information without damaging the broader structure of shared features, as excessive removal can harm retained-class accuracy.
More recently, studies have shown that neural representations are spatially and structurally localized Frantar and Alistarh (2023); Meng et al. (2022), motivating sparsity-based unlearning approaches that identify and mask a small set of influential weights to remove targeted information Fan et al. (2023); Jia et al. (2023); Pochinkov and Schoots (2024). Such methods have been successfully applied to both discriminative architectures Jia et al. (2023) and generative models Fan et al. (2023); Pochinkov and Schoots (2024) such as large language models, demonstrating that selective parameter masking can effectively achieve forgetting, while preserving global performance. Despite their efficiency, existing approximate unlearning approaches generally assume a centralized setup, where the model provider performs unlearning on a shared global model using the forget-set data. In contrast, our work considers unlearning in personalized models that have been locally adapted using private client data, aiming to remove provider-specified class information while retaining personalization-specific knowledge.
@@
ZKPs Goldreich and Oren (1994) enable verifiable ML by allowing a prover to demonstrate correct computation without revealing private inputs or intermediate states. ZK-SNARKs Bitansky et al. (2012); Kilian (1992); Micali (2000) are widely used for ML verification owing to their short proofs, fast verification, and modular circuits that can be partitioned across layers Maheri et al. (2025); Liu et al. (2021); Lee et al. (2024); Sun et al. (2024b); Zcash (2022). While suitable for large-scale auditing, proof generation remains costly in time and memory Kang et al. (2022).
Verification of machine unlearning ensures that models have genuinely forgotten target data. Existing methods include backdoor and sensitivity-based verification, which use poisoned or sensitivity of samples to test deletion Sommer et al. (2022); Gao et al. (2024); Guo et al. (2023); Zhou et al. (2025), but can be bypassed by a dishonest prover, and reproducing verification, which replays unlearning traces via proofs of learning Thudi et al. (2022b); Weng et al. (2024); Eisenhofer et al. (2025) but remains computationally heavy. Recent cryptographic frameworks instantiate proofs of unlearning with SNARKs and hash chains Eisenhofer et al. (2025), yet these verify exact unlearning (full retraining), making proof generation impractical. Furthermore, they are vulnerable to the stochasticity of SGD Zhang et al. (2024): an adversary can exploit minibatch randomness to retrain on retained samples whose gradients mimic those of the removed data, producing an apparently valid proof even though the model still encodes the forget-set information. This motivates ZK-friendly approximate unlearning methods that achieve verifiability without retraining.
@@ 2.1 Optimal Brain Surgeon
Second-order pruning methods such as Optimal Brain Damage LeCun et al. (1989) and Optimal Brain Surgeon Hassibi et al. (1993) introduced a principled approach to compensate for parameter removal by modeling loss curvature. By expanding the loss near an optimum, these methods derive closed-form updates that minimally affect performance after zeroing selected weights. Later work proposed scalable variants using empirical Fisher information Amari (1998), block-diagonal curvature Singh and Alistarh (2020), and structured sparsity Kurtic et al. (2022), enabling efficient use in pruning, quantization, and compression of modern networks Dong et al. (2017); Wang et al. (2019); Kuznedelev et al. (2024).
@@ 3 System and Problem Formulation
A client holds a private personalization set $D_{p}$ and obtains a personalized model by applying a fixed operator $P$ (e.g., short-horizon SGD, low-rank adapters such as LoRA Hu et al. (2022), or linear probes): $\theta_{p}\;=\;P(\theta_{0};\,D_{p}).$
@@
Parties and data. The model provider holds $\theta_{0}$, the pretraining corpus $D$, the forget/retain split $(D_{f},D_{r})$, and publishes traceability artifacts $\Psi$ (e.g., a sparse mask $m^{\star}$). The client holds private data $D_{p}$ and its personalized model $\theta_{p}=P(\theta_{0};D_{p})$, and executes $U(\cdot;\Psi)$ locally to obtain $\theta_{u}$. As in prior work Kang et al. (2022); Weng et al. (2023); Maheri et al. (2025), the model architecture is public; model weights ($\theta_{p},\theta_{u}$) remain private.
Workflow (reader map). At a high level, the provider publishes a public traceability artifact $\Psi$ that specifies the agreed unlearning transformation. Given $\Psi$, the client locally computes the unlearned model $\theta_{u}=U(\theta_{p};\Psi)$ and generates a zero-knowledge proof $\pi$ certifying operator compliance for the committed inputs and outputs. The provider then verifies $\pi$ against the public inputs $(\Psi,\operatorname{Com}(\theta_{p}),\operatorname{Com}(\theta_{u}))$, thereby certifying correct execution of $U$ without learning the client’s private model parameters or personalization data.
@@
Proof system primitives. In a ZK setting, verification is conducted through an arithmetic circuit that encodes the agreed computation $U$. A concise overview of circuit representations, constraint systems, and polynomial commitment schemes used in ZK-SNARKs is provided in Appendix B. Here, the client acts as the prover and generates a proof $\pi$ attesting that the committed personalized model $\theta_{p}$ was correctly transformed into $\theta_{u}$ under the public artifact $\Psi$. The provider acts as the verifier and checks the validity of $\pi$ without accessing any private inputs. The public inputs to the circuit are the traceability artifact $\Psi$ and the cryptographic commitments $\mathsf{Com}(\theta_{p})$ and $\mathsf{Com}(\theta_{u})$, which bind model parameters to their corresponding proofs without revealing them. This guarantee ensures that the agreed unlearning step was actually applied to the committed model and was not silently skipped or altered, thereby providing the basic trust anchor on which empirical forgetting evaluation builds. Formal definitions of the proof system properties—soundness, correctness, and zero-knowledge—as well as the binding and hiding guarantees of the commitment scheme are provided in Appendix C. We assume that a valid commitment $\mathsf{Com}(\theta_{p})$ has been established beforehand, for example through a proof-of-training or proof-of-personalization protocol Garg et al. (2023); Abbaszadeh et al. (2024); Sun et al. (2024a).
@@
Threat model. Adversaries are computationally bounded. (i) Privacy (curious verifier/provider): The verifier (including a curious provider) observes only the public inputs $(\Psi,\mathsf{Com}(\theta_{p}),\mathsf{Com}(\theta_{u}))$ and an argument $\pi$. By the zero-knowledge property (Appendix B) and the hiding property of $\mathsf{Com}(\cdot)$, the proof transcript reveals nothing about $\theta_{p}$, $\theta_{u}$, or $D_{p}$ beyond what is already implied by the public inputs. The model architecture is public, whereas all model weights and client-specific statistics remain part of the private witness. (ii) Integrity (dishonest client/prover): A dishonest prover may attempt to produce an accepting proof for some $\theta_{u}\neq U(\theta_{p};\Psi)$. By knowledge soundness, the verifier accepts only if the publicly committed $\theta_{u}$ equals the (deterministic) output of the agreed operator $U$ applied to the committed $\theta_{p}$ under $\Psi$. After acceptance, all subsequent predictions and updates must reference $\mathsf{Com}(\theta_{u})$, ensuring the deployed model is exactly the certified output of $U$. (iii) Black-box leakage (unlearning quality): An external party may have black-box access to the unlearned model outputs; this is not a protocol break. Any residual dependence on $D_{f}$ is treated as unlearning quality and evaluated separately in our experiments measured by MIA success in Table 1. Operationally, this means the next predictions are produced by the committed $\theta_{u}$, not by any model that still encodes forget-set patterns. All security properties are inherited from Halo2 Zcash (2022); formal definitions and assumptions appear in Appendix C.
@@
Scope. We do not address secure erasure of the client’s pre-unlearning model, as an archival copy of $\theta_{p}$ may persist on the device. Secure deletion is a challenging problem in its own right, often requiring trusted hardware-based method Hunt et al. (2018); Wu et al. (2024), and is considered orthogonal to our goal of verifiable unlearning. Our guarantee is verifiable use: after acceptance, any future inference or update must reference $\mathsf{Com}(\theta_{u})$, so deployed predictions are produced by the certified unlearned model, and do not rely on parameters that encode forget-set patterns. Operational controls (e.g., requiring a valid proof per update/inference) can enforce this usage constraint, but the physical deletion of prior weights is out of scope.
Procedural ZK guarantee vs. semantic unlearning.
An accepting proof certifies procedural correctness: the committed output satisfies $\theta_{u}=U(\theta_{p};\Psi)$ for the committed witness, while revealing neither $\theta_{p}$, $\theta_{u}$, nor $D_{p}$ beyond public inputs. Semantic forgetting is not a cryptographic guarantee; because $U$ is approximate relative to the gold standard in (2), unlearning quality is assessed empirically (e.g., via prediction-level alignment in (3)).
@@
Design goals for proof efficiency. We aim for ZK circuits that (i) minimize the number of constraints by verifying operator-specific identities instead of re-running heavy optimization (e.g., SGD) or training inside the circuit; (ii) utilise only linear operations (matrix–vector products, inner products), to avoid non-linear activations and large lookup tables that dominate SNARK cost Maheri et al. (2025); and (iii) avoid stochastic optimization inside the circuit — we verify deterministic operator identities and eliminate randomness  — thereby preventing prover adaptivity and the forging attacks observed in verification–unlearning settings that leverage SGD randomness Zhang et al. (2024).